### v1

# Correlation functions of harmonic lattices in d-dimensional space

Abstract
We study the correlation functions between the dynamical variables and between their conjugate momenta at sites of a harmonic lattice in the $d$-dimensional Euclidean space. We show that at the thermodynamic limit, they can be expressed in terms of Lauricella’s C-type hypergeometric series. Furthermore, using these expressions, we explicitly demonstrate that the correlators near the center of the lattice satisfying Diriclet boundary conditions coincide with those for the lattice with the periodic boundary conditions. By utilizing these expressions, we expect to make it easier to create programs that compute fast and highly precise for the quantum information quantities of subsystems within lattices.
1 Introduction
Indices of information loss due to entanglements, such as Renyi entropy are subjects of research not only in quantum information theory but also in many other fields, including field theory and condensed matter physics[1]. In the harmonic lattice model representing the discretized Klein Gordon field, to determine these quantities, one must calculate the vacuum correlators $Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ and $P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ for the set of sites of interest, where $q_{\vec{r}},q_{{\vec{s}}}$ and $p_{\vec{r}},p_{\vec{s}}$ are quantized dynamical variables of the sites with position vectors $\vec{r},\vec{s}$ and their conjugate momentum operators, respectively. Hereafter, we will simply refer to these correlators as $Q$ and $P$ or matrices $Q$,$P$ and so on as needed.
For example, consider an isolated lattice $U$ and a subset $A$ of sites, where we wish to obtain Renyi entropy of degree $n$, $S_{A}^{(n)}$, between A and its complement. The procedure is as follows: First, compute the matrices $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ restricted to sites $\vec{r}$ and $\vec{s}$ within $A$. Second, find the eigenvalues $\xi_{i}^{2}$ of $\Xi=QP$. The Renyi entropy can be computed using the following formula[2].
$S_{A}^{(n)}=\frac{1}{n-1}\sum_{i=1}^{\dim(\mathcal{H}_{A})}\log\left((\xi_{i}+\frac{1}{2})^{n}-(\xi_{i}-\frac{1}{2})^{n}\right),$ (1)
where $\mathcal{H}_{A}$ is the Hilbert space of the subsystem $A$ and $n\geq 2$ is an integer parameter.
For the computations on finite lattices to approximate those in the thermodynamic limit, the lattice spacing 1, the size $l_{A}$ of the subsystem $A$, and the entire lattice size $L$ must satisfy the constraint $1\ll l_{A}\ll L$. On the other hand, as $l_{A}$, $L$, and the spatial dimension $d$ increase, computation of the matrices $Q$ and $P$ becomes extremely time-consuming.
In this manuscript, we show that by considering the thermodynamic limit, the lattice scale $L\to\infty$ and replacing the sum over all site positions $(k_{1},k_{2},...,k_{d})$ within $U$ appearing in the computation of $Q$ and $P$ with multiple integrals $\int\int\dots\int dk_{1}dk_{2}\dots dk_{d}$ and calculating them analytically, the results for spatial dimensions of two or more are expressed by the C-type Lauricella’s hypergeometric series, $F_{C}$, irrespective of the boundary conditions for the dynamical variables $q_{r}$s and their conjugate momenta $p_{r}$s. We expect that by utilizing these expressions along with appropriate numerical computation libraries, it makes easier to create programs that rapidly compute various information quantities arising from entanglement.
2 Correlators on finite harmonic lattices
We consider a free scalar field in $(d+1)$ dimensional Lorentzian space-time and discretize it into a cubic lattice with $L^{d}$ sites of harmonic oscillators coupled with neighbor sites by spring-like interaction. The Hamiltonian is given by
$H=\frac{1}{2}\sum_{\vec{j}}(p_{\vec{j}}^{2}+\omega^{2}q_{\vec{j}}^{2}+\sum_{\alpha=1}^{d}(q_{\vec{j}+\vec{e}_{\alpha}}-q_{\vec{j}})^{2}),$ (2)
where $\vec{j}=(j_{1},j_{2},\ldots,j_{d})$ denotes the position of a site and $\vec{e}_{\alpha}=(e_{1}=0,0,\dots,e_{\alpha}=1,0,\ldots,0)$ is a unit vector on a spatial axis.
When the lattice $U$ satisfies periodic boundary conditions(PBCs) along directions of spatial axes, $q_{\vec{j}+L\vec{e}_{\alpha}}=q_{\vec{j}},\ p_{\vec{j}+L\vec{e}_{\alpha}}=p_{\vec{j}},\ \alpha=1,2,\ldots,d$, based on the variables $q_{\vec{j}}$s and $p_{\vec{j}}$s obtained by solving the Hamiltonian equations, the dispersion relation $\omega_{k}$ and $Q$, $P$ matrices are given [4] by
$\omega_{\vec{k}}^{(P)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}(\pi k_{\alpha}/L)}\ ,$ (3)
$Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\frac{1}{\omega_{\vec{k}}^{(P)}}\prod_{\alpha=1}^{d}\cos[2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})/L],$ (4)
$P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\omega_{\vec{k}}^{(P)}\prod_{\alpha=1}^{d}\cos[2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})/L],$ (5)
where $k_{\alpha}$’s are integers such that $0\leq k_{\alpha}\leq L-1$ and $V=L^{d}$.
In the case where U satisfies Dirichlet boundary conditions(DBCs) as follows:
$q_{(j_{1},j_{2},\ldots,j_{\alpha}=L,\ldots,j_{d})}$ $=q_{(j_{1},j_{2},\ldots,\ j_{\alpha}=0,\ \ldots,j_{d})}=0,$
$\ p_{(j_{1},j_{2},\ldots,j_{\alpha}=L,\ldots,j_{d})}$ $=p_{(j_{1},j_{2},\ldots,\ j_{\alpha}=0\ ,\ldots,j_{d})}=0,$
the dispersion relation and $Q$, $P$ matrices are given [4] by:
$\omega_{\vec{k}}^{(D)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}(\pi k_{\alpha}/(2L))},$ (6)
$Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\frac{1}{\omega_{\vec{k}}^{(D)}}\prod_{\alpha=1}^{d}2\sin\left(\frac{\pi r_{\alpha}k_{\alpha}}{L}\right)\sin\left(\frac{\pi s_{\alpha}k_{\alpha}}{L}\right),$ (7)
$P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\omega_{\vec{k}}^{(D)}\prod_{\alpha=1}^{d}2\sin\left(\frac{\pi r_{\alpha}k_{\alpha}}{L}\right)\sin\left(\frac{\pi s_{\alpha}k_{\alpha}}{L}\right),$ (8)
where, as in the case of PBCs, $V=L^{d}$. Note that, as long as the size $L$ of $U$ is finite, for any spatial axis direction, $j_{\alpha}$ ranges 0 to $L$ and the number of $j_{\alpha}$s is $L+1$, but truly dynamical $q_{(j_{1},j_{2},\ldots,j_{\alpha},\ldots,j_{d})}$ and $p_{(j_{1},j_{2},\ldots,j_{\alpha},\ldots,j_{d})}$ exist only at $1\leq j_{\alpha}\leq L-1$ and $k_{\alpha}$s also take integers in the range $1\leq k_{\alpha}\leq L-1$.
3 Correlation functions in the thermodynamic limit
In this section, we show that the correlation functions $Q_{\vec{r}\vec{s}}$ in (4),(7) and $P_{\vec{r}\vec{s}}$ in (5), (8) are both expressed by the same type of hypergeometric series in the thermodynamic limit, regardless of the boundary conditions. In order to discuss common properties beyond differences in $Q$ and $P$, we also denote $Q_{\vec{r}\vec{s}}$ as $G^{(1/2)}_{\vec{r}\vec{s}}$, $P_{\vec{r}\vec{s}}$ as $G^{(-1/2)}_{\vec{r}\vec{s}}$ for the case that the entire lattice satisfies PBCs and $J_{\vec{r}\vec{s}}^{(\pm 1/2)}$ for the lattice satisfying DBCs.
First, in correlators in the finite lattice satisfying PBCs, (4) and (5), taking the limit as $L\to\infty$ , and replacing $2\pi k_{\alpha}/L\to k_{\alpha}$, $\sum_{\vec{k}}\to\int_{0}^{2\pi}\dots\int_{0}^{2\pi}L^{d}/(2\pi)^{d}d^{d}k$ yields the following expression:
$G^{(\pm 1/2)}_{\vec{r}\vec{s}}=\frac{1}{2(2\pi)^{d}}\int_{0}^{2\pi}\dots\int_{0}^{2\pi}d^{d}k\ (\omega_{k}^{(P)})^{\mp 1}\prod_{\alpha=1}^{d}\cos n_{\alpha}k_{\alpha},$ (9)
where $n_{\alpha}$ denotes $|r_{\alpha}-s_{\alpha}|$. The dispersion relation and its reciprocal are expressed by
$(\omega_{k}^{(P)})^{\pm 1}=(\omega^{2}+2\sum_{\alpha=1}^{d}(1-\cos k_{\alpha}))^{\pm 1/2}=(\Omega)^{\pm 1}(1-v\sum_{\alpha=1}^{d}\cos k_{\alpha})^{\pm 1/2}$
$=\Omega^{\pm 1}\sum_{m=0}^{\infty}\begin{pmatrix}\pm 1/2\\ m\end{pmatrix}(-v)^{m}\left(\sum_{\alpha}^{d}\cos k_{\alpha}\right)^{m}=\Omega^{\pm 1}\sum_{m=0}^{\infty}\frac{(\mp 1/2)_{m}}{m!}v^{m}\left(\sum_{\alpha=1}^{d}\cos k_{\alpha}\right)^{m},$ (10)
[…]
$G^{(\pm 1/2)}_{\vec{r}\vec{s}}=\frac{\Omega^{\mp 1}}{2(2\pi)^{d}}\sum_{m=0}^{\infty}\frac{(\pm 1/2)_{m}}{m!}v^{m}\underbrace{\int_{0}^{2\pi}\dots\int_{0}^{2\pi}d^{d}k\ \left(\sum_{\alpha=1}^{d}\cos k_{\alpha}\right)^{m}\prod_{\alpha^{\prime}=1}^{d}\cos n_{\alpha^{\prime}}k_{\alpha^{\prime}}}_{I_{m,n_{1},n_{2},\ldots,n_{d}}},$ (11)
The multiple integral $I_{m,n_{1},n_{2},\ldots,n_{d}}$ in (11) can be expressed as a linear combination of the products of integrals with respect to each $k_{\alpha}$ by expanding $(\sum_{\alpha=1}^{d}\cos k_{\alpha})^{m}$ as follows:
$I_{m,n_{1},\dots,n_{d}}=\sum_{l_{1}=0}^{m}\sum_{l_{2}=0}^{m}\dots\sum_{l_{d}=0}^{m}\delta_{m,(\sum_{\alpha^{\prime}=1}^{d}l_{\alpha^{\prime}})}\frac{m!}{l_{1}!l_{2}!\cdots l_{d-1}!l_{d}!}\left(\prod_{\alpha=1}^{d}I_{l_{\alpha},n_{\alpha}}\right),$ (12)
where $I_{l_{\alpha},n_{\alpha}}$ is given by
$I_{l_{\alpha},n_{\alpha}}=\int_{0}^{2\pi}\cos^{l_{\alpha}}k_{\alpha}\cos n_{\alpha}k_{\alpha}dk_{\alpha}=\begin{cases}\frac{2\pi}{2^{l_{\alpha}}}\begin{pmatrix}l_{\alpha}\\ (l_{\alpha}-n_{\alpha})/2\end{pmatrix}&\text{if}\ l_{\alpha}\geq n_{\alpha}\ \text{and}\ l_{\alpha}=n_{\alpha}\hskip-5.69054pt\mod 2,\\ 0&\text{otherwise}.\end{cases}$ (13)
Putting $(l_{\alpha}-n_{\alpha})/2\overset{\underset{\mathrm{def}}{}}{=}p_{\alpha}$, we have:
$\frac{m!}{l_{1}!l_{2}!\cdots l_{d-1}!l_{d}!}\prod_{\alpha=1}^{d}\begin{pmatrix}l_{\alpha}\\ (l_{\alpha}-n_{\alpha})/2\end{pmatrix}=\frac{m!}{\prod_{\alpha}p_{\alpha}!(n_{\alpha}+p_{\alpha})!}=\frac{m!}{\prod_{\alpha}p_{\alpha}!n_{\alpha}!(n_{\alpha}+1)_{p_{\alpha}}}.$ (14)
Furthermore, when we put $N\overset{\underset{\mathrm{def}}{}}{=}n_{1}+n_{2}+\dots+n_{d}$, $p_{1}+p_{2}+\dots+p_{d}=(m-N)/2\overset{\underset{\mathrm{def}}{}}{=}u$ holds. Let $a=\pm 1/2$, then we obtain the following equation:
$\left(\pm\frac{1}{2}\right)_{m}=(a)_{N+2u}=(a)_{N}(a+N)_{2u}=(a)_{N}2^{2u}\left(\frac{a+N}{2}\right)_{u}\left(\frac{a+1+N}{2}\right)_{u}.$ (15)
Using (12),(13),(14) and (15), the equation (11) can be written as
$G^{(\pm 1/2)}_{\vec{r}\vec{s}}=\frac{\Omega^{\mp 1}}{2}\frac{(a)_{N}}{\prod_{\alpha}n_{\alpha}!}\frac{v^{N}}{2^{N}}\sum_{p_{1}=0}^{\infty}\sum_{p_{2}=0}^{\infty}\dots\sum_{p_{d}=0}^{\infty}\frac{\left(\frac{a+N}{2}\right)_{u}\left(\frac{a+1+N}{2}\right)_{u}}{\prod_{\alpha}(n_{\alpha}+1)_{p_{\alpha}}}\frac{v^{2u}}{\prod_{\alpha}p_{\alpha}!}\ .$ (16)
The infinite series appearing in equation (16) are Lauricella’s C-type hypergeometric series defined by:
$F^{(d)}_{C}(\alpha,\beta;\gamma_{1},\gamma_{2},\dots,\gamma_{d};z_{1},z_{2},\dots,z_{d})=\sum_{p_{1}=0}^{\infty}\sum_{p_{2}=0}^{\infty}\dots\sum_{p_{d}=0}^{\infty}\frac{(\alpha)_{u}(\beta)_{u}}{\prod_{\alpha=1}^{d}(\gamma_{\alpha})_{p_{\alpha}}}\frac{\prod_{\alpha=1}^{d}z_{\alpha}^{p_{\alpha}}}{\prod_{\alpha=1}^{d}p_{\alpha}!}\ ,$ (17)
and the $Q$ and $P$ correlation functions are expressed, respectively, as follows:
$Q_{\vec{r}\vec{s}}^{(P)}$ $=\frac{1}{2\Omega}\frac{\left(\frac{1}{2}\right)_{N}}{\prod_{\alpha}n_{\alpha}!}\frac{v^{N}}{2^{N}}F_{C}^{(d)}\left(\frac{N+1/2}{2},\frac{N+3/2}{2};n_{1}+1,n_{2}+1,\ldots,n_{d}+1;v^{2},v^{2},\ldots,v^{2}\right),$ (18)
$P_{\vec{r}\vec{s}}^{(P)}$ $=\frac{\Omega}{2}\frac{\left(-\frac{1}{2}\right)_{N}}{\prod_{\alpha}n_{\alpha}!}\frac{v^{N}}{2^{N}}F_{C}^{(d)}\left(\frac{N-1/2}{2},\frac{N+1/2}{2};n_{1}+1,n_{2}+1,\ldots,n_{d}+1;v^{2},v^{2},\ldots,v^{2}\right).$ (19)
The convergence condition for $F_{C}^{(d)}(\alpha,\beta;\gamma_{1},\ldots,\gamma_{d};v^{2},v^{2},\dots,v^{2}$) appearing in the expressions (18) and (19) is $v\cdot d<1$ or $v\cdot d=1\wedge\ \gamma_{1}+\gamma_{2}+\dots+\gamma_{d}-\alpha-\beta>0$. Since $v=2/(\omega^{2}+2d)$, in the massive case, the former condition is always satisfied. Even in the massless case, $\omega^{2}=0$, since $(n_{1}+1)+\dots+(n_{d}+1)-(a+N)/2-(a+1+N)/2=d-a-1/2$, the latter convergence condition is satisfied for $d\geq 2$.
Next, we consider the thermodynamic limit of the correlators, (7) and (8) in the lattice under DBCs. Replacing $\pi k_{\alpha}/L\to k_{\alpha}$, $\sum_{\vec{k}}\to\int_{0}^{\pi}\dots\int_{0}^{\pi}L^{d}/\pi^{d}d^{d}k$ yields the following expression:
$J_{\vec{r}\vec{s}}^{(\pm 1/2)}=\frac{1}{2\pi^{d}}\int_{0}^{\pi}\dots\int_{0}^{\pi}d^{d}k(\omega_{k}^{(D)})^{\mp 1}\prod_{\alpha=1}^{d}(\cos n_{0}^{(\alpha)}k_{\alpha}-\cos n_{1}^{(\alpha)}k_{\alpha}),\\$ (20)
where
$n_{\epsilon_{\alpha}}^{(\alpha)}=\begin{cases}|r_{\alpha}-s_{\alpha}|,\ \text{if\ }\epsilon_{\alpha}=0,\\ r_{\alpha}+s_{\alpha},\ \text{if\ }\epsilon_{\alpha}=1.\end{cases}$ (21)
Hereafter, unless confusion arises, the superscript $(\alpha)$ will be omitted. The dispersion relation is given by the following equation, as in (10):
$\omega_{k}^{(D)}=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}k_{\alpha}/2}=\Omega\sqrt{1-v\sum_{\alpha=1}^{d}\cos k_{\alpha}},\text{\ \ \ where\ \ \ }\Omega^{2}=\omega^{2}+2d,\ v=\frac{2}{\Omega^{2}}.$ (22)
Expanding $\prod_{\alpha=1}^{d}(\cos n_{0}^{(\alpha)}k_{\alpha}-\cos n_{1}^{(\alpha)}k_{\alpha})$ yields the following expression:
$J_{\vec{r}\vec{s}}^{(\pm 1/2)}$ $=\sum_{\epsilon_{1}=0}^{1}\sum_{\epsilon_{2}=0}^{1}\dots\sum_{\epsilon_{d}=0}^{1}(-1)^{\sum_{\alpha=1}^{d}\epsilon_{\alpha}}J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}$, (23)
$J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}=\frac{1}{2\pi^{d}}\int_{0}^{\pi}\dots\int_{0}^{\pi}d^{d}k(\omega_{k}^{(D)})^{\mp 1}\prod_{\alpha=1}^{d}\cos n_{\epsilon_{\alpha}}k_{\alpha}.$ (24)
The difference between (9) and (24), besides $n_{\alpha}$ changing to $n_{\epsilon_{\alpha}}$, are that (i)the coefficient before the integral changes from $1/(2\cdot 2^{d}\pi^{d})$ to $1/(2\pi^{d})$ and that (ii) the integration range for each $k_{\alpha}$ changes from $0,2\pi]$ to $0,\pi]$. Among these, due to (ii), the value of the integral for each $k_{\alpha}$ in (24) becomes half that of (13), and the entire multiple integral becomes multiplied by $1/2^{d}$. This cancels exactly with the change due to (i), so (24) matches (9) and its calculated result (16) except for the substitution from $n_{\alpha}$ to $n_{\epsilon_{\alpha}}$:
$J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}=\frac{\Omega^{\mp 1}}{2}\frac{(a)_{N}}{\prod_{\alpha}n_{\epsilon_{\alpha}}!}\frac{v^{N}}{2^{N}}\sum_{p_{1}=0}^{\infty}\sum_{p_{2}=0}^{\infty}\dots\sum_{p_{d}=0}^{\infty}\frac{\left(\frac{a+N}{2}\right)_{u}\left(\frac{a+1+N}{2}\right)_{u}}{\prod_{\alpha}(n_{\epsilon_{\alpha}}+1)_{p_{\alpha}}}\frac{v^{2u}}{\prod_{\alpha}p_{\alpha}!},$ (25)
where we set $N\overset{\underset{\mathrm{def}}{}}{=}n_{\epsilon_{1}}+n_{\epsilon_{2}}+\dots+n_{\epsilon_{d}}$, $u\overset{\underset{\mathrm{def}}{}}{=}p_{1}+p_{2}+\dots p_{d}$ as before. Then correlation functions $Q$ and $P$ under DBCs can be expressed, respectively, using Lauricella’s hypergeometric series as follows:
$\hskip-8.5359ptQ_{\vec{r}\vec{s}}^{(D)}$
$\hskip-8.5359pt=\sum_{\epsilon_{1}=0}^{1}\dots\sum_{\epsilon_{d}=0}^{1}(-1)^{\sum_{\alpha}\epsilon_{\alpha}}\frac{1}{2\Omega}\frac{\left(\frac{1}{2}\right)_{N}}{\prod_{\alpha}n_{\epsilon_{\alpha}}!}\frac{v^{N}}{2^{N}}F_{C}^{(d)}\left(\frac{N+1/2}{2},\frac{N+3/2}{2};n_{\epsilon_{1}}+1,\ldots,n_{\epsilon_{d}}+1;v^{2},\dots,v^{2}\right),$ (26)
$\hskip-8.5359ptP_{\vec{r}\vec{s}}^{(D)}$
$\hskip-8.5359pt=\sum_{\epsilon_{1}=0}^{1}\dots\sum_{\epsilon_{d}=0}^{1}(-1)^{\sum_{\alpha}\epsilon_{\alpha}}\frac{\Omega}{2}\frac{\left(-\frac{1}{2}\right)_{N}}{\prod_{\alpha}n_{\epsilon_{\alpha}}!}\frac{v^{N}}{2^{N}}F_{C}^{(d)}\left(\frac{N-1/2}{2},\frac{N+1/2}{2};n_{\epsilon_{1}}+1,\dots,n_{\epsilon_{d}}+1;v^{2},\dots,v^{2}\right).$ (27)
Thus, we have shown that for harmonic lattices in space with $d\geq 2$, the correlation functions for the dynamical variables in the thermodynamic limit are expressed as a hypergeometric series of the same type, regardless of the boundary conditions of the lattice.
Reference [3] derives the necessary $Q$ and $P$ matrices in the thermodynamic limit to discuss the contour functions for the entanglement entropies in one-dimensional harmonic lattice. By setting $d=1$ in the $Q$ and $P$ matrices given in (18), (19) and (26), (27), and performing some calculations, we obtain all the correlation functions in the appendix C of [3].
4 Connection between correlators of different boundary conditions
In the thermodynamic limit, near the center of the lattice, $r_{\alpha}\sim s_{\alpha}\sim L/2$, the physical states are expected to be not largely affected by boundary conditions, and the correlation functions under the DBCs (26) and (27) should coincide with those under PBCs (18) and (19). On the other hand, for $J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}$ of (25), when all $n_{\epsilon_{\alpha}}=n_{0}=|r_{\alpha}-s_{\alpha}|$, it coincides precisely with the PBC correlators (16). Therefore, we examine the behavior of the other terms appearing in (23), (26) and (27) in the thermodynamic limit, i.e, the behavior of terms $J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}$ with some $n_{\epsilon_{\alpha}}=n_{1}=r_{\alpha}+s_{\alpha}\sim L\to\infty$. Since $J^{(\pm 1/2)}_{n_{\epsilon_{1}},n_{\epsilon_{2}},\dots,n_{\epsilon_{d}}}$ is symmetric under permutations of $n_{\epsilon_{\alpha}}$, setting $\epsilon_{1}=\epsilon_{2}=\dots=\epsilon_{d_{1}}=1$, $\epsilon_{d_{1}+1}=\dots=\epsilon_{d}=0$, we do not lose generality. So, we set $n_{1}^{(1)}+n_{1}^{(2)}+...+n_{1}^{(d_{1})}\overset{\underset{\mathrm{def}}{}}{=}N_{+}$ and $N_{-}\overset{\underset{\mathrm{def}}{}}{=}N-N_{+}$. Hereafter in this section, $n_{\epsilon_{\alpha}}^{(\alpha)}$ in formulae will be denoted by $n_{\alpha}$ for simplicity.
We evaluate the behavior of each factor in (25) as $n_{1}^{(\alpha)}\to\infty$. Using Stirling’s formula, we obtain
$(a)_{N}$ $\sim(N_{+}+N_{-}+a)^{N_{+}+N_{-}+a-1/2}e^{-(N_{+}+N_{-}+a)}\sim N_{+}^{N_{+}}\left(1+\frac{N_{-}+a}{N_{+}}\right)^{N_{+}}N_{+}^{N_{-}+a-1/2}e^{-N_{+}}$
$\sim N_{+}^{N_{+}}e^{N_{-}}e^{-N_{+}}N_{+}^{N_{-}+a-1/2}.$ (28)
For $\alpha=1,2,\dots,d_{1}$, let $n_{\epsilon_{\alpha}}^{(\alpha)}=\lambda_{\alpha}N_{+}$, we obtain the following expression:
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim(\prod_{\alpha=1}^{d_{1}}n_{\alpha}^{n_{\alpha}+1/2}e^{-N_{+}})^{-1}\sim(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (29)
The formula for the maximum Shannon entropy of a information source with $d_{1}$ variables whose probability distribution is $(\lambda_{1},\lambda_{2},\dots,\lambda_{d_{1}})$ yields the following inequality:
$\log(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}=N_{+}\sum_{\alpha=1}^{d_{1}}(-\lambda_{\alpha}\log\lambda_{\alpha})\leq N_{+}\log d_{1}=\log d_{1}^{N_{+}}.$
This applies to (29) and we obtain
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim d_{1}^{N_{+}}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (30)
Using (28) and (30) and $v=2/(\omega^{2}+2d)$, the evaluation of the coefficient of the infinite series in (25) is as follows:
$\frac{\Omega^{\mp 1}}{2}\frac{(a)_{N}}{\prod_{\alpha}n_{\alpha}!}\frac{v^{N}}{2^{N}}\sim\left(\frac{d_{1}/d}{2}\right)^{N_{+}}O(N_{+}^{N_{-}+a-(1+d_{1})/2}).$ (31)
Evaluation of one term in the hypergeometric series yields the following expression:
$\frac{\left(\frac{a+N}{2}\right)_{u}\left(\frac{a+1+N}{2}\right)_{u}}{\prod_{\alpha}(n_{\alpha}+1)_{p_{\alpha}}}\frac{v^{2u}}{\prod_{\alpha}p_{\alpha}!}\sim\frac{(N_{+}/2)^{2u}}{(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{p_{\alpha}})N_{+}^{u}}\frac{v^{2u}}{\prod_{\alpha=1}^{d_{1}}p_{\alpha}!}\sim O((N_{+}v^{2})^{u})\prod_{\alpha=1}^{d_{1}}(1/\lambda_{\alpha})^{p_{\alpha}}.$ (32)
When some $\epsilon_{\alpha}$s are equal to $1$, the behavior of each term in (25) as $n_{1}^{(\alpha)}\sim N_{+}\to\infty$ is obtained by multiplying equations (31) and (32) as follows:
$\left(\frac{d_{1}}{2d}\right)^{N_{+}}O(N_{+}^{N_{-}+a-(1+d_{1})/2})O((N_{+}v^{2})^{u})\prod_{\alpha=1}^{d_{1}}(1/\lambda_{\alpha})^{p_{\alpha}}.$ (33)
Since each $(p_{1},p_{2},\dots,p_{d})$ term of infinite series in (25) is obtained in the thermodynamic limit ($n^{(\alpha)}_{1}\sim L\to\infty$), $p_{\alpha}$s in the term should be regarded as finite values. So the absolute value of (33) decreases exponentially and becomes zero. Thus, only $J_{n_{0}n_{0},\ldots,n_{0}}$ survives on the right-hand side of (23), and the correlation functions for the DBC case $J_{\vec{r}\vec{s}}^{(\pm 1/2)}$ coincide with that for the PBC case, namely $G_{\vec{r}\vec{s}}^{(\pm 1/2)}$ in (16).
5 Conclusions and future tasks
For harmonic lattices, we have shown that the vacuum correlation function matrices $Q$ and $P$ for the dynamical variables and their conjugate momenta on sites are consistently expressed using Lauricella’s C-type hypergeometric series $F_{C}^{(d)}$ in the thermodynamic limit, in any spatial dimension $d$ of the lattice. For lattices satisfying periodic boundary conditions, $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ are given by (18) and (19), and for those satisfying the Dirichlet boundary conditions, they are given by (26) and (27). These hypergeometric series absolutely converge for spatial dimensions $d\geq 2$. Additionally, we have specifically observed that the correlation functions $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ for sites near the center of the lattice under Dirichlet boundary conditions coincide with those in the lattice under periodic boundary conditions. In the specific cases, it is known that Lauricella’s C-type hypergeometric series is factorized into the product of Gauss’s hypergeometric series $_{2}F_{1}$s. As an example, let us consider $F_{C}^{(2)}$ appearing in the $Q$ of (18) for $d=2$ which is also also referred as Appell’s $F_{4}$ hypergeometric series. For the massless harmonic lattice, the variables of this hypergeometric series are $v^{2}=1/4$ and it is factorized as follows[5]:
$F_{4}\left(\frac{N+1/2}{2},\frac{N+3/2}{2};n_{1}+1,n_{2}+1;1/4,1/4\right)$
$={}_{2}F_{1}\left(\frac{N+1/2}{2},\frac{N+3/2}{2};n_{1}+1;1/2\right){}_{2}F_{1}\left(\frac{N+1/2}{2},\frac{N+3/2}{2};n_{2}+1;1/2\right).$ (34)
Further investigation into whether this expression is written in a more concise form, and whether any of the other correlation functions can be factorized in this manner, remains a future task.
Numerical calculations of $Q$ and $P$ matrices are essential for research using harmonic lattices on entanglement indices like Renyi entropy and their corner functions, edge terms and contour functions, which represents the contribution to the indices from vertices, edges on the boundary of the physical system and from each site within it. However, as the dimensionality increases, creating programs to perform these computations quickly under memory constraints for sufficiently large-scale subsystems becomes increasingly difficult. The analytical solutions presented here, which express the correlation functions in $d$-dimensional space using hypergeometric series, are expected to facilitate the creation of these programs.
The following two issues remain: 1.We will investigate whether (34) and other hypergeometric series in correlator functions can be written in more concise forms. 2.For $d=2$ and $d=3$, we will obtain corner functions and edge terms for specific shapes and scales of subsystems based on the $Q$ and $P$ values computed by the program cord that employs expressions (18) and (19). Then, we will compare these results with those from previous studies, which were computed based on the $Q$ and $P$ values in the finite-scale lattice of (4) and (5) or on those of integral expressions in (11).